OpenStudy (anonymous):

factor the difference of two squares x^2-4

OpenStudy (accessdenied):

Difference of squares: (a^2 - b^2) = (a + b)(a - b)

OpenStudy (anonymous):

(x+2)(x-2) x=2,-2